# View-Consistent Hierarchical 3D Segmentation Using Ultrametric Feature Fields

### v1

@@ 1 Introduction
Unlike prior work, we explicitly aim for our scene segmentation to be hierarchical. This means that the scene segmentation has a tree structure, where the root group is the full scene, and any group can be recursively divided into smaller groups, all the way down to the point level. A group is defined as a spatially connected neighborhood where all pairwise feature distances are within some threshold. At first glance, this is a familiar contrastive learning problem: within-segment feature pairs should have small distances, and cross-segment feature pairs should have large distances, so that thresholding yields groups that follow segmentation boundaries. However, as we demonstrate in our experiments, this is not sufficient to create a consistent multi-level hierarchy. We find it is crucial to use ultrametric distances, rather than standard Euclidean distances, in the contrastive loss. In an ultrametric space, for any three points , , and , distances satisfy a condition stronger than the standard triangle inequality , namely that . Ultrametric spaces are ideally suited for hierarchical clustering because distance-based groupings are transitive: if and , then it follows that . In other words, if we group and together and and together, then we automatically also group and together. Thus, as grows, smaller clusters are naturally merged into larger clusters, automatically giving rise to a hierarchy.
@@
We evaluate our approach on models from the PartNet dataset, showing that our method recovers a view-consistent hierarchy of segmentations that captures the natural part decomposition of a 3D object. Furthermore, we introduce a synthetic dataset with hierarchical segmentation annotations based on the NeRF Blender Dataset. Unlike PartNet, our proposed dataset offers hierarchical decomposition of more complex scenes. In all evaluations, we compare our method to an array of competitive baselines, measuring IoU accuracy and 3D consistency, and we demonstrate that our method outperforms existing open-vocabulary 3D segmentation methods, such as DFF [24], LeRF [19], and SAM-3D [6]. Additionally, we introduce a metric for measuring the quality of a hierarchy, and demonstrate that the segmentations in prior work lack hierarchical structure, while our output is hierarchical by construction.
@@ 2.0.1 Neural Radiance and Feature Fields
One line of research aimed to learn a semantic or instance branch alongside the NeRF’s color and density, thus yielding view-consistent segmentation [60, 30, 48, 25, 43]. Semantic-NeRF [60] was a seminal work that learned a semantic field to propagate 2D segmentations into new views. Instance-NeRF [30] improved Semantic-NeRF by handling panoptic segmentation from Mask2Former [11]. Nevertheless, all these methods are restricted to a closed vocabulary of instance categories, and additionally do not attempt any hierarchical understanding of the segmented instances.
Another recent direction has been to learn a generic volumetric feature field alongside NeRF. DFF [24] and LeRF [19] distilled 2D image features generated by off-the-shelf feature extractors such as CLIP [39], DINO [5], and LSeg [26] into feature fields that enable 3D segmentation and editing in NeRF. However, these methods often struggle to recover precise segmentation boundaries, and neither establishes a hierarchical structure on the features. In contrast, we learn a feature field that distills a hierarchy of segmentations from noisy multi-view SAM predictions.
Yin et al. [56] and Cen et al. [6] used NeRF to propagate a SAM query into novel views and establish a view-consistent segmentation. However, these methods can only handle one segmentation at a time, while our approach trains a feature field to jointly aggregate and reconcile hundreds of noisy 2D masks.
@@ 2.0.2 Hierarchical Segmentation
Hierarchical segmentation, a specialized domain within image segmentation, partitions an image into regions that exhibit a hiearchical tree structure, (i.e., each segmentation can be recursively divided into smaller segmentations). In the pre-deep learning era, researchers utilized non-parametric approaches to generate contours and enable hierarchical segmentation [44, 4, 36, 1, 2]. Ultrametric distance and the watershed transform were employed to identify hierarchical clusters based on RGB values [36, 4, 1]. Building upon these foundations, subsequent advancements by Yarkony et al. [55] and Xu et al. [53] improved the efficiency and flexibility of the ultrametric hierarchical segmentation method. More recently, rather than focusing on the hierarchical structure within the image, Zhao et al. [59] and Li et al. [27] shifted their focus to segmentation within the context of a class hierarchy. The concept of class hierarchy has also been explored in the field of human parsing, where segmentation is performed according to the hierarchical structure of the human body [13, 50, 51, 15]. Mo et al. [34] introduced PartNet, a large scale 3D mesh dataset annotated with hierarchical segmentation. Some researchers have focused on exploring this dataset’s hierarchical structure and produce multi-granularity segmentation on point clouds, as demonstrated by Sun et al. [46] and Sun et al. [45].
@@ 3.0.3 Segment Anything Model (SAM)
SAM [22] is a state-of-the-art vision foundation model for image segmentation. Given a query in the form of points, a mask, a box, or a language prompt, SAM predicts a segmentation that best reflects the prompt. SAM predicts segmentations at three levels of granularity: instance, part, and subpart.
While SAM produces excellent results on single images, it is nontrivial to lift its predictions into 3D. This is because SAM’s predictions are not consistent across viewpoints. View-dependent effects such as occlusion, lighting, camera zoom, and camera position can influence the regions that SAM segments, and also influence which “granularity" SAM assigns. For example, a pen may be segmented into multiple parts in a close-up view, then segmented as a single instance in a wider view, and then be completely missed when viewed from farther away. Furthermore, it is unclear how to best lift outputs from a “queryable" design into a well-organized 3D representation: given an abundance of queries, SAM will generate an abundance of masks, which overlap with one another unpredictably, and have no straightforward unification.
@@ 4 Method
The previous section described the core components of our approach. In this section we describe how these pieces fit together, and explain the training and inference pipeline.
@@ 4.1.2 Contrastive Learning
Minimizing the loss in Eq. 3 across many pairs and yields a feature space where segmentations can be recovered by querying the neighborhood around a point, and taking the region where pairwise feature distances are all below some threshold. As discussed in earlier sections, whether or not the resulting segmentation will be hierarchical depends on whether the distance metric is an ultrametric or a standard Euclidean metric. Our main goal is to reduce ultrametric distances, but to improve optimization, we apply loss in both ultrametric space and Euclidean space:
@@
As described in Eqs. 1 and 2, ultrametric distances are defined on a graph. At each training step, we form an approximate graph by sampling 4096 pixels within the image, and connect each pixel to its 10 nearest neighbors, and use feature distances as edge weights. Minimizing the ultrametric distance of a positive pair means finding the shortest path between the vertices, and penalizing the longest edge on that path. We use the binary partition tree algorithm [12] to efficiently compute ultrametric distances during training.
@@ 4.1.3 Hierarchical Sampling
Because our input segmentations are 2D, overlapping, and multi-granular, it is ambiguous whether a pair of pixels “lie within the same mask” and therefore represent a positive pair in our contrastive formulation. For example, two pixels may be within the same mask at a coarse granularity but not for a finer granularity. We address this ambiguity with a simple tree-based strategy for sampling positive and negative pairs of pixels, which is presented in the supplementary.
For each image, we organize the segmentation masks into a hierarchical structure determined by the inclusion ratio between them. One mask is designated as a child of another mask when and . We empirically set and . We additionally include an all-positive mask as the root of the tree.
For each view, we sample positive and negative pairs starting from the leaf nodes of the hierarchical tree, i.e. masks at the finest granularity. For a leaf mask with parent mask , we randomly select two pixels in and designate these as a positive pair. We then randomly select one pixel in and one pixel in and designate these as a negative pair. We then move to the parent mask, and repeat.
Since the masks grow as we rise up the hierarchy, pairs which are declared negative at the bottom will eventually be declared positive higher up. Our approach mitigates this conflict by computing the contrastive loss within each level and summing the result across levels. This asks that the positive distances be smaller than the negative distances within a single level, and avoids the possibility of a pair being simultaneously positive and negative in Eq. 3. In all, our sampling strategy reflects an ultrametric structure and helps the segmentation hierarchy propagate across the scene.
@@ 4.1.4 Improving Depth with Segmentation
In our approach, the 3D structure of the scene helps resolve conflicts in segmentation cues. Can segmentation cues also help us resolve ambiguity in 3D structure? To explore this possibility, we add an assumption that regions belonging to the same semantic mask have smoothly changing depth. We propose a regularization term that penalizes changes in curvature (i.e. the third derivative of depth) within a segment. Concretely, at each training iteration, we sample local patches of pixels, ensuring that each sampled patch lies within one of the SAM’s finest-grained mask predictions. The depth continuity loss is defined as
Here, represents the ray angle difference between adjacent pixels, denotes a threshold, and consists all sampled pixels for depth continuity loss, with denoting adjacent pixels in a row or column. To mitigate the impact of the depth continuity loss on density convergence, we only apply this loss from the midpoint of the training process. Fig. 5 presents a visual comparison of depth estimation with and without the depth continuity loss.
@@ 4.2.2 Segmenting in 3D
To achieve 3D-consistent segmentation, we begin by creating a featurized 3D point cloud, using the feature maps and depth maps obtained from our trained implicit field. Following related NeRF works, we remove outliers and downsample the pointcloud using Open3D [61]. We create a graph from the points, denoted as , where includes all points in the point cloud, and consists of edges connecting points with their nearest neighbors determined by 3D distance, and feature distances provide edge weights. Given a threshold indicating the level of granularity, we eliminate all edges longer than the threshold. The remaining connected components in the graph represent segments at the specified level of granularity. Since this process occasionally yields tiny unusable segments, we retain the largest components as our final segmentation. Using different values of allows for segmentation at varying levels of granularity.
To transfer our 3D segmentation to a novel view, we begin by estimating depth for the view using the implicit field, which allows us to connect the novel image to the graph. Then, for each pixel, we find its nearest neighbors within a distance threshold in the graph, and assign a segmentation label using the mode of the neighborhood.
@@ 5 Experiments
We report qualitative and quantitative evaluations for our method and a set of relevant baselines. In Sec. 5.1 we give an overview of the datasets we evaluate on, and we define our evaluation metrics. In Sec. 5.2 and Sec. 5.3, we present quantitative comparisons and ablation studies. We additionally visualize our method’s hierarchical and view-consistent segmentation on real-world scenes from the Tanks and Temples Dataset [23]. We present implementation details in Sec. 5.4.
@@ 5.1 Datasets and Metrics
The official PartNet dataset [34, 7] contains 3D models with professionally verified hierarchical part decomposition. Furthermore, each object model is rendered into 24 viewpoints that uniformly cover the viewing sphere. PartNet offers us the unique ability to evaluate on hierarchical segmentation verifiably aligns with human perception of structural hierarchy. We evaluate on five objects from each of the chair, table, and storage furniture categories, totalling to 15 objects models.
@@
While PartNet contains 3D objects with hierarchical part annotations, each model is visually simplistic and lacks the photometric complexity often associated with NeRFs. To the best of our knowledge, no publicly available dataset exists that displays both complex photometric structure as well as 3D-consistent hierarchical segmentation labels. Thus, we create a new synthetic dataset based on the Blender Dataset [31] which we call Blender with Hierarchical Segmentation (Blender-HS). We choose three scenes from the Blender Dataset that exhibit clear hierarchical structure: Lego, Hotdog, and Drums, and for each we define three granularity levels, which we denote as “scene", “collection", and “object". The object level consists of each distinct 3D asset; the “collection” level contains 3D asset groupings that were decided by the original artist; the “scene” level consists of the entire scene. While many potential hierarchies exist in general, the artist-defined groupings naturally represent a hierarchy that makes sense for humans. We present visualizations of the hierarchical ground truth in the supplementary.
@@
In hierarchical segmentation, it is important that each pixel belongs to only one mask for each level of granularity. To measure this, we propose a metric which we call Segmentation Injectivity (SI). For each ground truth segmentation, we retrieve its best-matching predicted mask and sample two random points and from that mask. We then query the model for a new mask at each of these points, at the same granularity. We compute the Intersection over Union (IoU) between the two resulting masks. Note that a perfect model will return the same mask at both locations, since these points belong to the same ground-truth segment, whereas a worse model will return masks which may not even overlap. Since the random point sampling introduces some randomness, we run this 100 times per ground truth mask, and average the scores across all ground truth masks and all viewpoints. We find this is more tractable than estimating masks densely, and in practice exhibits low variance. Note there is a trade-off between the SI score and the NC score: predicting an excessive number of overlapping masks could inflate the NC score but would significantly lower the SI score.
@@
We use a View Consistency (VC) score to measure the 3D consistency of image segmentations. The key idea of this score is to estimate segmentations in two nearby viewpoints independently, and then warp these estimates onto one another, and measure their agreement. To compute this score, we begin by defining a pixel transformation from a source viewpoint to another one shifted 10 degrees, using ground truth depth and camera pose. Next, for each ground truth segmentation, we retrieve its best-matching predicted mask, sample a random point from that mask, and compute the point’s location in the shifted viewpoint . We then obtain segmentation estimates to compare, centered on in the first view and centered on in the second. Finally we warp ’s mask into the viewpoint, and measure IoU with the mask. To ensure that the IoU does not merely reveal occlusion/disocclusion differences, we use ground truth visibility masks to remove pixels which are invisible in either view, before computing the IoU. Similar to the other scores, we mitigate noise induced by the random sampling by computing this score 100 times per ground truth mask, and average the scores across all ground truth masks and all viewpoints.

### v2

@@ 1 Introduction
Unlike prior work, we explicitly aim for our scene segmentation to be hierarchical. This means that the scene segmentation has a tree structure, where the root group is the full scene, and any group can be recursively divided into smaller groups, all the way down to the point level. A group is defined as a spatially connected neighborhood where all pairwise feature distances are within some threshold. At first glance, this is a familiar contrastive learning problem: within-segment feature pairs should have small distances, and cross-segment feature pairs should have large distances, so that thresholding yields groups that follow segmentation boundaries. However, we demonstrate this typical setup is not sufficient to create a consistent hierarchy, and we find it is crucial to use ultrametric distances, rather than standard Euclidean distances, in the contrastive loss. In an ultrametric space, for any three points , , and , distances satisfy a condition stronger than the standard triangle inequality , namely that . Ultrametric spaces are ideally suited for hierarchical clustering because distance-based groupings are transitive: if and , then it follows that . In other words, if we group and together and and together, then we automatically also group and together. Thus, as grows, smaller clusters are naturally merged into larger clusters, automatically giving rise to a hierarchy.
@@
We evaluate our approach on models from the PartNet dataset, showing that our method recovers a view-consistent hierarchy of segmentations that captures the natural part decomposition of a 3D object. Furthermore, we introduce a synthetic dataset with hierarchical segmentation annotations based on the NeRF Blender Dataset. Unlike PartNet, our proposed dataset offers hierarchical decomposition of more complex scenes. In all evaluations, we compare our method to a set of competitive baselines, measuring IoU accuracy and 3D consistency, and we demonstrate that our method outperforms existing open-vocabulary 3D segmentation methods, such as DFF [26], LeRF [21], and SAM-3D [7]. Additionally, we introduce a metric for measuring the quality of a hierarchy, and demonstrate that the segmentations in prior work lack hierarchical structure, while our output is hierarchical by construction.
@@ 2.0.1 Neural Radiance and Feature Fields
One line of research aimed to learn a semantic or instance branch alongside the NeRF’s color and density, thus yielding view-consistent segmentation [63, 32, 51, 27, 45]. Semantic-NeRF [63] was a seminal work that learned a semantic field to propagate 2D segmentations into new views. Instance-NeRF [32] improved Semantic-NeRF by handling panoptic segmentation from Mask2Former [12]. Nevertheless, these methods are restricted to a closed vocabulary of instance categories, and additionally do not attempt a hierarchical understanding of the segmented instances.
Another recent direction has been to learn a generic volumetric feature field alongside NeRF. DFF [26], N3F [49], and LeRF [21] distilled 2D image features generated by off-the-shelf feature extractors such as CLIP [41], DINO [5], and LSeg [28] into feature fields that enable 3D segmentation and editing in NeRF. However, these methods often struggle to recover precise segmentation boundaries, and also do not establish a hierarchical structure on the features. In contrast, we learn a feature field that distills a hierarchy of segmentations from noisy multi-view SAM predictions.
Finally, other works [59, 7, 6] used NeRF or 3D Gaussians [20] to propagate a single SAM query into novel views and establish view-consistency. However, these methods handle one segmentation at a time, while our approach trains a feature field to jointly aggregate and reconcile hundreds of noisy 2D masks.
@@ 2.0.2 Hierarchical Segmentation
Hierarchical segmentation, a specialized domain within image segmentation, partitions an image into regions that exhibit a hiearchical tree structure, (i.e., each segmentation can be recursively divided into smaller segmentations). In the pre-deep learning era, researchers utilized non-parametric approaches to generate contours and enable hierarchical segmentation [46, 4, 38, 1, 2]. Ultrametric distance and the watershed transform were employed to identify hierarchical clusters based on RGB values [38, 4, 1], and subsequent advancements by Yarkony et al. [58] and Xu et al. [56] improved the efficiency and flexibility of ultrametric hierarchical segmentation methods. More recently, Zhao et al. [62] and Li et al. [29] shifted their focus from hierarchical segmentation within an image toward estimating the hierarchy of segmentation classes. This class hierarchy has also been of particular interest in the field of human parsing, where segmentation is performed according to the hierarchical structure of the human body [14, 53, 54, 16]. In the 3D domain, Mo et al. [36] introduced PartNet, a large scale 3D mesh dataset annotated with hierarchical segmentation, and researchers have subsequently explored hierarchical structure in 3D shapes, including multi-granularity segmentation on point clouds [48, 47, 65].
@@ 3.0.3 Segment Anything Model (SAM)
SAM [24] is a state-of-the-art vision foundation model for image segmentation. Given a query in the form of points, a mask, a box, or a language prompt, SAM predicts a segmentation that best reflects the prompt. SAM generates segmentations at three levels of granularity: instance, part, and subpart.
While SAM produces excellent results on single images, it is nontrivial to lift its predictions into 3D. This is because SAM’s predictions are not consistent across viewpoints. For example, a pen may be segmented into multiple parts in a close-up view, then segmented as a single instance in a wider view, and then be completely missed when viewed from farther away. Furthermore, it is unclear how to best lift outputs from a “queryable" design into a well-organized 3D representation: given an abundance of queries, SAM will generate an abundance of masks, which overlap with one another unpredictably, and have no straightforward unification.
@@ 4 Method
The previous section described the core components of our approach. In this section we describe how these pieces fit together, and explain the training and inference pipelines.
@@ 4.1.2 Contrastive Learning
Minimizing the loss in Eq. 3 across many pairs and yields a feature space where segmentations can be recovered by querying a point and finding the neighborhood where pairwise feature distances are all below some threshold. As discussed in earlier sections, whether or not the resulting segmentation will be hierarchical depends on whether the distance metric is an ultrametric or a standard Euclidean metric. Our main goal is to reduce ultrametric distances, but to improve optimization, we apply loss in both ultrametric space and Euclidean space:
@@
As described in Eqs. 1 and 2, ultrametric distances are defined on a graph. At each training step, we form an approximate graph by sampling 4096 pixels within the image, connect each pixel to its 10 nearest neighbors, and use feature distances as edge weights. We use the binary partition tree algorithm [13] to efficiently compute ultrametric distances during training.
@@ 4.1.3 Hierarchical Sampling
Because our input segmentations overlap within and across views, it is ambiguous whether a pair of pixels “lie within the same mask” and therefore represent a positive pair in our contrastive formulation. For example, two pixels may be within the same mask at a coarse granularity but not for a finer granularity. We address this ambiguity with a simple tree-based strategy for sampling positive and negative pairs of pixels.
For each image, we organize the segmentation masks into a hierarchical structure determined by the inclusion ratio between them. We additionally include an all-positive mask as the root of the tree. Then, for each view, we sample positive and negative pairs starting from the leaf nodes of the hierarchical tree, i.e. masks at the finest granularity. For a leaf mask with parent mask , we randomly select two pixels in and designate these as a positive pair. We then randomly select one pixel in and one pixel in and designate these as a negative pair. We then move to the parent mask and repeat. Additional details are provided in the supplementary.
We note that masks grow in size as we rise up the hierarchy, and pairs which were declared negative at finer granularity will be declared positive in courser segmentations. We mitigate this conflict by computing the contrastive loss within each level. This asks that the positive distances be smaller than the negative distances within a single level, and avoids the possibility of a pair being simultaneously positive and negative in Eq. 3. Taken together, our sampling strategy reflects an ultrametric structure and helps the segmentation hierarchy propagate across the scene.
@@ 4.1.4 Improving Depth with Segmentation
Our approach shows that the 3D structure of a scene can resolve conflicts in segmentation cues. Conversely, can segmentation cues help resolve ambiguity in 3D structure? To explore this possibility, we add an assumption that regions belonging to the same semantic mask have smoothly changing depth. We propose a regularization that penalizes changes in curvature (i.e. the third derivative of depth) within a segment. Concretely, at each training iteration, we sample local patches of pixels, ensuring that each sampled patch lies within one of the SAM’s finest-grained mask predictions. The depth continuity loss is defined as
where represents the ray angle difference between adjacent pixels, denotes a threshold, and contains the sampled pixels, with denoting adjacent pixels in a row or column. For training stability, we start using this loss halfway through training. Fig. 5 presents the benefits of the depth continuity loss.
@@ 4.2.2 Segmenting in 3D
To achieve 3D-consistent segmentation, we create a featurized 3D point cloud by unprojecting feature maps and depth maps rendered from the optimized implicit field. Following related NeRF works, we remove outliers and downsample the pointcloud using Open3D [64]. Then, we construct a k-nearest-neighbor (KNN) graph on the 3D point cloud, and we set each edge weight in the KNN graph to the feature distance between the connected vertices. For a threshold indicating the level of granularity, we remove all edges longer than that threshold, and the remaining connected components represent segments at the specified level of granularity. In practice, we keep the largest components as our final segmentation. Using different values of allows for segmentation at varying levels of granularity.
To propagate the 3D segmentation into a novel view, we first render a depth map of the novel view and unproject the render into 3D points. Then, for each unprojected pixel, we find its nearest neighbors to the previously estimated KNN graph and assign a segmentation label using the mode of the neighborhood.
@@ 5 Experiments
We report qualitative and quantitative evaluations for our method and a set of relevant baselines. In Sec. 5.1 we give an overview of the datasets we evaluate on, and we define our evaluation metrics. In Sec. 5.2 and Sec. 5.3, we present quantitative comparisons and ablation studies. We present implementation details in Sec. 5.4.
@@ 5.1 Datasets and Metrics
The official PartNet dataset [36, 8] contains 3D models with professionally verified hierarchical part decomposition. Furthermore, each object model is rendered into 24 viewpoints that uniformly cover the viewing sphere. PartNet offers us the unique ability to evaluate on hierarchical segmentation that verifiably aligns with human perception of structural hierarchy. We evaluate on five objects from each of the chair, table, and storage furniture categories, totalling to 15 objects models.
@@
While PartNet contains 3D objects with hierarchical part annotations, each model is visually simplistic and lacks the photometric complexity often associated with NeRFs. To the best of our knowledge, no publicly available dataset exists that displays both complex photometric structure as well as 3D-consistent hierarchical segmentation labels. Thus, we create a new synthetic dataset based on the Blender Dataset [33] which we call Blender with Hierarchical Segmentation (Blender-HS). We choose three scenes from the Blender Dataset that exhibit clear hierarchical structure: Lego, Hotdog, and Drums, and for each we define three granularity levels, which we denote as “scene”, “collection”, and “object”. The object level consists of each distinct 3D asset; the “collection” level contains 3D asset groupings that were decided by the original artist; the “scene” level consists of the entire scene. While many potential hierarchies exist in general, the artist-defined groupings naturally represent a hierarchy that makes sense for humans. We present visualizations of the hierarchical ground truth in the supplementary. We use 100 views in the training set for training and 10 views from the validation set for evaluation.
@@
In hierarchical segmentation, it is important that each pixel belongs to only one mask for each level of granularity. To measure this, we propose a metric which we call Segmentation Injectivity (SI). For each ground truth segmentation, we retrieve its best-matching predicted mask and sample two random points and from that mask. We then query the model for a new mask at each of these points, at the same granularity. We compute the Intersection over Union (IoU) between the two resulting masks. Note that a perfect model will return the same mask at both locations, since these points belong to the same ground-truth segment, whereas a worse model will return masks which may not even overlap. Since the random point sampling introduces randomness, we run this 100 times per ground truth mask and average the scores across all ground truth masks and all viewpoints. We find this is more tractable than estimating masks densely, and in practice exhibits low variance. Note there is a trade-off between the SI score and the NC score: predicting an excessive number of overlapping masks could inflate the NC score but would significantly lower the SI score.
@@
We use a View Consistency (VC) score to measure the 3D consistency of image segmentations. The key idea of this score is to estimate segmentations in two nearby viewpoints independently, and then warp these estimates onto one another, and measure their agreement. To compute this score, we begin by defining a pixel transformation from a source viewpoint to another one shifted 10 degrees, using ground truth depth and camera pose. Next, for each ground truth segmentation, we retrieve its best-matching predicted mask, sample a random point from that mask, and compute the point’s location in the shifted viewpoint . We then obtain segmentation estimates to compare, centered on in the first view and centered on in the second. Finally we warp ’s mask into the viewpoint, and measure IoU with the mask. To ensure that the IoU does not merely reveal occlusion/disocclusion differences, we use ground truth visibility masks to remove pixels which are invisible in either view, before computing the IoU. Similar to the other scores, we compute this score 100 times per ground truth mask and average the scores across all ground truth masks and all viewpoints.